OpenStudy (anonymous):

If i is imaginary what is i^20?

OpenStudy (rulnick):

1

OpenStudy (rulnick):

... i^2 = -1, i^4 = 1.

OpenStudy (anonymous):

Thanks so much everyone!!!!

OpenStudy (rulnick):

welcome

OpenStudy (anonymous):

i^2 = -1 therefor: (i^20) = (i^2)^10 = (-1)^10 = 1

OpenStudy (amistre64):

\[(i^2)^{10}\]maybe